# On admissibility in post-hoc hypothesis testing

### v2

@@ 2.1 Admissible test families
In the standard hypothesis testing paradigm we search for tests which have the lowest type-II error (highest power) among all those which have type-I error bounded by $\alpha$ (level-$\alpha$ tests). A level-$\alpha$ test is better than another to the extent that it has higher power, and a test is uniformly most powerful (UMP) if it has the highest power among all level-$\alpha$ tests. In our framework, we compare two test families that are type-I risk safe by examining their losses under the alternative. In particular, the analogy of a UMP test is an “admissible” test family.
@@
We call $\phi$ strictly preferable to $\delta$ if the inequality is strict for at least one $B\in\Gamma$ and $Q\in\mathcal{Q}$. If there exists some type-I risk safe $\phi$ that is strictly preferable to $\delta$ relatively to $\Gamma$ we say that $\delta$ is $\Gamma$-inadmissible. If no decision family $\phi$ is strictly preferable to $\delta$ relative to $\Gamma$ then we say that $\delta$ is $\Gamma$-admissible. (Said differently, $\delta$ is $\Gamma$-admissible if the following holds: whenever there exists a $\phi$ that is weakly preferable to $\delta$, (2.5) actually holds with equality for all $B\in\Gamma$ and $Q\in\mathcal{Q}$.)
We leave implicit in Definition 2.4 whether we’re considering binary or randomized tests; the definition holds for both. We will always make clear in the text which kinds of tests are under consideration.
In what follows we will always assume that any collection $\Gamma$ of adversaries under consideration includes the constant mappings. This will simplify many of the proofs. It is also intuitive: it would be odd if an adversary could play $L_{b}(1,0)$ for only a subset of the observations. To formalize this assumption, we will consider only those $\Gamma\supseteq\mathbf{C}$ where
@@
In the sequel, we will often shorthand $B(X)$ to $B$ in equations to save space. Note that for any $b\in\mathcal{B}$, $L_{b}(1,\delta(X,b))=L_{b}(1,0)(1-\delta(X,b))$ (recall we are assuming that $L_{b}(1,1)=L_{b}(0,0)=0$), so (2.5) can be equivalently written as $\mathbb{E}_{X\sim Q}L_{B}(1,0)\delta(X,B)\leq\mathbb{E}_{X\sim Q}L_{B}(1,0)\phi(X,B)$. In particular, this implies that if $\phi(X,b)\geq\delta(X,b)$ for some $b\in\mathcal{B}$ and all $X\in F$ for some $F\subseteq\mathcal{X}$, then $\mathbb{E}_{Q}[L_{b}(1,\phi(X,b))\mathbf{1}\left\{X\in F\right\}]\leq\mathbb{E}_{Q}[L_{b}(1,\delta(X,b)\mathbf{1}\left\{X\in F\right\}]$, a fact we will use often.
It is worth spending a moment to relate Definition 2.4 to a more familiar setting. Suppose there is just a single loss function $L_{b}$. Then (2.5) is equivalent to $\mathbb{E}_{Q}[\delta(X,b)]\leq\mathbb{E}_{Q}[\phi(X,b)]$. Thus $\phi$ is strictly preferable to $\delta$ (relative to $\mathbf{U}$, which is a singleton if there is one loss function) if it is more powerful than $\delta$, meaning it rejects with higher probability under the alternative. Moreover, per the discussion after Definition 2.3, if $\alpha=L^{-1}(0,1)$, then a type-I risk safe test has type-I error bounded by $\alpha$. We may therefore conclude:
Observation 2.5.
For a single loss $L$ (the traditional Neyman-Pearson setting) with $\alpha=L^{-1}(0,1)$, any admissible test family is a uniformly most powerful level-$\alpha$ test.
One could of course consider other notions of admissibility. Grünwald (2024) considers a family $\phi$ to be preferable to $\delta$ if, for any adversary, $\phi$ rejects whenever $\delta$ rejects with probability 1. We compare our notion of admissibility with that of Grünwald (2024) further in Appendix B, but suffice it to say here that requiring that $\phi$ be better than $\delta$ with probability 1 is a strong requirement, and does not recover the notion of a uniformly most powerful test.
@@
Remark 2.6.
Unless we consider $\Gamma=\mathbf{U}$ there is an asymmetry between the adversaries considered in the definition of type-I risk and the definition of $\Gamma$-admissibility. Lemma B.1 shows that a type-I risk safe rule has expected type-I loss bounded by 1 across any mapping $B:\mathcal{X}\to\mathcal{B}$, while in $\Gamma$-admissibility we are deliberately restricting the mappings under consideration, but only for the alternative (i.e. for type II risk). This is by design. Our first goal is to find tests which have bounded type-I risk under even the worst case adversary, as this seems the most appropriate goal for a post-hoc hypothesis test. Only then do we consider type-II risk. This is analogous to how Neyman and Pearson prioritize type-I error bounded by $\alpha$ and only then look for tests with higher power.
@@
    $\Gamma=\mathbf{U}$. The first natural set of adversaries to consider is all mappings $B:\mathcal{X}\to\mathcal{B}$. Under such a liberal notion of admissibility, a great many test families are admissible. Indeed, under mild assumptions on the losses, every test family $\delta$ which satisfies $\mathcal{R}_{\mathcal{P}}(\delta)=1$ is admissible. In other words, every test which is defined by a sharp e-value is admissible. This is somewhat unsatisfying, and leads us to consider the following more restricted class of adversaries.
@@
    $\Gamma=\mathbf{C}$ for $\mathbf{C}$ as in (2.6). We will discuss the interpretation of $\mathbf{C}$ further in Section 5.3, but let us mention here that it may helpful to interpret $\mathbf{C}$-admissibility in terms of counterfactuals. Namely, if $\delta$ is $\mathbf{C}$ admissible then for every type-I risk safe $\phi$, there exists some loss $b$ such that if the adversary had played $L_{b}$, we would not regret playing $\delta$ instead of $\phi$. We will give several necessary conditions for a randomized test family to be $\mathbf{C}$-admissible and provide a full characterization of $\mathbf{C}$-admissible binary test families.
Note that $\mathbf{U}$ and $\mathbf{C}$ are, respectively, the largest and smallest sets of adversaries we can consider under the assumption that $\Gamma\supseteq\mathbf{C}$: we are studying the two extrema of Definition 2.4. Table 2 gives an overview of our admissibility results.
@@
That is, letting the losses take on any value in $1,\infty)$ lets us consider any post-hoc significance level $\alpha\in(0,1)$. Arguably, the right hand side of (2.7) is what one would expect of a definition of post-hoc hypothesis testing (see, e.g., Wang and Ramdas (2022, Section 6.1), Koning (2023) and Gauthier et al. (2025)).
@@ 2.2 Notation, assumptions, and further background
We will also assume that there exists some $b$ such that $L_{b}(0,1)>1$. Otherwise a test family may always reject with probability 1 and remain type-I risk safe, making the problem uninteresting. This assumption is akin to not studying $\alpha=1$ in the traditional Neyman-Pearson paradigm, which is uninteresting for the same reason. We also make the necessary (quite modest) measure-theoretic assumptions so that (2.3) holds. See Appendix B.1 for precise details.
Finally, going forward, we assume that we are testing a point null $\mathcal{P}=\{P\}$ against a point alternative $\mathcal{Q}=\{Q\}$.
@@ 3 Warmup: Likelihood ratio tests
In the case of a single loss function, the Neyman-Pearson lemma (Neyman and Pearson, 1933) implies that the likelihood ratio test is uniformly most powerful. What can we say about likelihood ratio-style tests in a setting with multiple losses?
@@
$P(\phi^{\textsc{np}}(X,b)=1)=L_{b}^{-1}(0,1).$
We use the superscript “NP” to refer to Neyman-Pearson. Note that the second term $\gamma\mathbf{1}\left\{\Lambda(X)=\kappa(b)\right\}$ is only relevant for discrete distributions. We refer to Shao (2008) for a modern treatment of the Neyman-Pearson lemma and the optimality of the likelihood ratio test.
Suppose that a test family $\delta$ plays as $\phi^{\textsc{np}}(\cdot,b^{*})$ on some $b^{*}$. How does it behave on $b\neq b^{*}$? Naively, one would like to play $\phi^{\textsc{np}}(\cdot,b)$ on each $b$. Such a test family would not be type-I risk safe, however. This is easy to see, even in the case of two losses. Consider playing $\phi^{\textsc{np}}(X,b_{1})$ on $L_{b_{1}}$ and $\phi^{\textsc{np}}(X,b_{2})$ on $L_{b_{2}}$ with $b_{2}>b_{1}$. Since $L_{b_{2}}(0,1)>L_{b_{1}}(0,1)$ we have $\kappa({b_{2}})>\kappa({b_{1}})$. If an adversary $B$ plays loss $b_{2}$ for $X$ with $\Lambda(X)\geq\kappa({b_{2}})$ and $b_{1}$ for $X$ with $\kappa({b_{1}})\leq\Lambda(X)<\kappa({b_{2}})$, then $\mathbb{E}_{P}[L_{B}(0,\delta(X,B))]=L_{b_{2}}(0,1)P(\Lambda(X)\geq\kappa({b_{2}}))+L_{b_{1}}(0,1)P(\kappa(b_{1})\leq\Lambda(X)<\kappa({b_{2}}))>1$, so $\mathcal{R}_{P}(\delta)>1$.
However, for $b\leq b^{*}$, $\delta(\cdot,b)$ can play as $\phi^{\textsc{np}}(\cdot,b^{*})$ and remain type-I risk safe. In fact, in the discrete case, for those $X$ such that $\Lambda(X)=\kappa(b)$ it can even raise the probability of rejection due to the extra freedom afforded by the fact that $L_{b}(0,1)\leq L_{b^{*}}(0,1)$. For $b>b^{*}$ meanwhile, $\delta(\cdot,b)$ must lower the probability of rejection, but may still reject. Proposition 3.1 demonstrates that, for any $\Gamma\supseteq\mathbf{C}$, the only admissible test family which plays as $\phi^{\textsc{np}}(\cdot,b^{*})$ on $b^{*}$ has the following form:
@@
This representation is perhaps more intuitive to digest. If an adversary $B$ plays $B(X)=b^{*}$ for all $X$, the expected type-I risk $\mathbb{E}_{P}[L_{B}(1,\delta(X,B)]$ is precisely 1 by definition of the likelihood ratio test. For $b\leq b^{*}$ we have $L_{b}(0,1)\leq L_{b^{*}}(0,1)$ so an adversary playing such a $b$ can only lower the type-I risk and $\delta$ can play as $\phi^{\textsc{np}}(X,b^{*})$. For $b>b^{*}$, $\delta$ may still reject, but it must lower its rejection probability to make up for the increase in type-I risk. We set $\delta$ such that $L_{b}(0,1)\delta(X,b)=L_{b^{*}}(0,1)\phi^{\textsc{np}}(X,b^{*})$ so that the type-I risk stays the same.

### v3

@@ 2.1 Admissible test families
In the standard theory of hypothesis testing, two classes of tests are of interest: uniformly most powerful (UMP) tests and admissible tests. A level-$\alpha$ test (i.e., having type-I error bounded by $\alpha$) is UMP if it has greater or equal power than every other level-$\alpha$ test for every $Q\in\mathcal{Q}$. If $\mathcal{Q}$ is composite, then UMP tests only exist in special settings. These include practically relevant cases, however, such as the one-sided t-test and $\chi^{2}$-tests, and more generally when $\mathcal{P}\cup\mathcal{Q}$ admit a monotone likelihood ratio. An admissible test, meanwhile, cannot be uniformly improved across all $Q\in\mathcal{Q}$ (meaning that it cannot be strictly improved for some $Q$ and not made worse on any other $Q$). In this work we generalize admissibility to a post-hoc setting.
@@
We call $\phi$ strictly preferable to $\delta$ if the inequality is strict for at least one $B\in\Gamma$ and $Q\in\mathcal{Q}$. If there exists some type-I risk safe $\phi$ that is strictly preferable to $\delta$ relatively to $\Gamma$ we say that $\delta$ is $\Gamma$-inadmissible. If no decision family $\phi$ is strictly preferable to $\delta$ relative to $\Gamma$ then we say that $\delta$ is $\Gamma$-admissible. (Here is an alternative definition: $\delta$ is $\Gamma$-admissible if for every $\phi$ that is weakly preferable to $\delta$, (2.5) actually holds with equality for all $B\in\Gamma$ and $Q\in\mathcal{Q}$.)
Remark 2.5.
Unless we consider $\Gamma=\mathbf{U}$ there is an asymmetry between the adversaries considered in the definition of type-I risk and the definition of $\Gamma$-admissibility. Lemma B.2 shows that a type-I risk safe rule has expected type-I loss bounded by 1 across any mapping $B:\mathcal{X}\to\mathcal{B}$, while in $\Gamma$-admissibility we are deliberately restricting the mappings under consideration, but only for the alternative (i.e. for type II risk). This is by design. In the worst case, the loss may be data-dependent—we want to ensure we remain type-I risk safe in this case. However, we might also be cautiously optimistic that it won’t be data-dependent, and allow ourselves to choose a test that is $C$-admissible based on this hope, all the while knowing that if we are wrong then at least we are protected under the null. Therefore, our first goal is to find tests which have bounded type-I risk under even the worst case adversary. Only then do we consider type-II risk. This is analogous to how Neyman and Pearson prioritize type-I error bounded by $\alpha$ and only then look for tests with higher power.
We leave implicit in Definition 2.4 whether we are considering binary or randomized tests; the definition holds for both. We will always make clear in the text which kinds of tests are under consideration.
In what follows we will always assume that any collection $\Gamma$ of adversaries includes the constant mappings. This will simplify many of the proofs. It is also intuitive: it would be odd if an adversary could play $L_{b}(1,0)$ for only a subset of the observations. To formalize this assumption, we will consider only those $\Gamma\supseteq\mathbf{C}$ where
@@
In the sequel, we will often shorthand $B(X)$ to $B$ in equations to save space. Recall that for any $b\in\mathcal{B}$, $L_{b}(1,\delta(X,b))=L_{b}(1,0)(1-\delta(X,b))$ (we are assuming that $L_{b}(1,1)=L_{b}(0,0)=0$), so (2.5) can be equivalently written as $\mathbb{E}_{X\sim Q}L_{B}(1,0)\delta(X,B)\leq\mathbb{E}_{X\sim Q}L_{B}(1,0)\phi(X,B)$. In particular, this implies that if $\phi(X,b)\geq\delta(X,b)$ for some $b\in\mathcal{B}$ and all $X\in F$ for some $F\subseteq\mathcal{X}$, then $\mathbb{E}_{Q}[L_{b}(1,\phi(X,b))\mathbf{1}\left\{X\in F\right\}]\leq\mathbb{E}_{Q}[L_{b}(1,\delta(X,b)\mathbf{1}\left\{X\in F\right\}]$, a fact we will use often.
Despite the differences between being UMP and admissible, it’s worth investigating when the two notions coincide. Suppose that there is only one loss function $L_{b_{0}}$. If $\phi$ is UMP at level $L_{b_{0}}^{-1}(0,1)$ then it satisfies $\mathbb{E}_{Q}[\delta(X,b_{0})]\leq\mathbb{E}_{Q}[\phi(X,b_{0})]$ for all $Q\in\mathcal{Q}$ and all other tests $\delta$. Thus, in particular, these inequalities are satisfied when $\delta$ is admissible (which notion of admissibility does not matter, since there is only one loss), thus implying that the inequalities must be equalities. Hence $\delta$ has the same power as $\phi$ on each $Q$, and is also UMP. We may therefore conclude:
Observation 2.6.
For a single loss $L$ (the traditional Neyman-Pearson setting) with $\alpha=L^{-1}(0,1)$, if a uniformly most powerful level-$\alpha$ test exists then any admissible test family is such a test.
One could consider other notions of admissibility. grunwald2024beyond considers a family $\phi$ to be preferable to $\delta$ if, for any adversary, $\phi$ rejects whenever $\delta$ rejects with probability 1. We compare our notion of admissibility with that of grunwald2024beyond further in Appendix B, but suffice it to say here that requiring that $\phi$ be better than $\delta$ with probability 1 is a strong requirement, and does not recover the notion of a uniformly most powerful test.
@@
    $\Gamma=\mathbf{U}$. The first natural set of adversaries to consider is all mappings $B:\mathcal{X}\to\mathcal{B}$. Under such a liberal notion of admissibility, a great many test families are admissible. Indeed, under mild assumptions on the losses, every test family $\delta$ which satisfies $\mathcal{R}_{\mathcal{P}}(\delta)=1$ is admissible. In other words, every test which is defined by a sharp e-value is admissible.
@@
    $\Gamma=\mathbf{C}$ for $\mathbf{C}$ as in (2.6). We discuss the interpretation of $\mathbf{C}$ further in Section 5.3, and provide a concrete motivation in Example 2.8, but let us mention here that it may be helpful to interpret $\mathbf{C}$-admissibility in terms of counterfactuals. Namely, if $\delta$ is $\mathbf{C}$-admissible then for every type-I risk safe $\phi$, there exists some loss $b$ such that if the adversary had played $L_{b}$, we would not regret playing $\delta$ instead of $\phi$. We will give several necessary conditions for a randomized test family to be $\mathbf{C}$-admissible and provide a full characterization of $\mathbf{C}$-admissible binary test families.
Note that $\mathbf{U}$ and $\mathbf{C}$ are, respectively, the largest and smallest sets of adversaries we can consider under the assumption that $\Gamma\supseteq\mathbf{C}$: we are studying the two extrema of Definition 2.4. While we do not explicitly study other sets of adversaries here (apart from when our results easy generalize to handle all families $\Gamma$; cf. Section 4), we believe that other sets could be of interest, as illustrated by Example 2.8 below, which also provides further motivation for studying $\mathbf{C}$-admissibility.
Table 2 gives an overview of our admissibility results.
@@
That is, letting the losses take on any value in $1,\infty)$ lets us consider any post-hoc significance level $\alpha\in(0,1)$. Arguably, the right hand side of (2.7) is what one would expect of a definition of post-hoc hypothesis testing (see, e.g., wang2022false, koning2023markov and gauthier2025values).
Example 2.8 (Example 1.1, Continued).
Consider the following idealized instantiation(s) of Example 1.1. Suppose $b$ represents the cost of investment expressed in some monetary unit; for concreteness say the cost is $\mathdollar b\cdot 10^{5}$. A funder is given the opportunity to invest $b$ in the drug’s production.
If the drug turns out to be ineffective yet an investment is made then the investor loses their capital. We formalize this as $L_{b}(0,0)=0,L_{b}(0,1)=b$. When the drug is effective, we consider two pay-off variations: 1. The pay-off is a fixed constant $C^{*}$. This reflects, for example, a situation in which the funder is the sole investor, and the investment is used to build a factory that can produce a fixed quantity of the drug. We view the loss as the regret incurred if the funder does not invest in the successful drug, so $L_{b}(1,1)=0$ and $L_{b}(1,0)=C^{*}$ for all $b$. 2. The pay-off is a strictly increasing function $f(b)$ of $b$. This could for example be the case if larger $b$’s represent a larger factory (more productive capacity) or if there are many funders who each get a share of profits proportional to their investment (in that case $f$ would be linear). In this case, again viewing $L_{b}(1,\cdot)$ as the regret of not investing in a successful drug, we set $L_{b}(1,0)=f(b)$ and $L_{b}(1,1)=0$. In the first variation, $L_{b}(1,0)=C^{*}$ is independent of $b$, showing that $\Gamma=\mathbf{C}$ is an important special case to consider. In the second variation, $L_{b}(1,0)$ is increasing in $b$. It then seems reasonable to assume that the investment $b$ is increasing in the amount of evidence for effectiveness—in practice the ‘adversaries’ might be the data-analysts in a company, who suggest to management to make a larger investment the more evidence they see in the data. This reflects how p-hacking typically occurs in practice: an analyst is tempted to lower $\alpha$ (increase $L_{b}(0,1)$, which, in the second case also increases $L_{b}(1,0)$) after seeing promising results. This suggests an analysis in which $\Gamma$ is restricted to those $B(X)$ which are (monotonically) increasing in a sufficient statistic of the data (e.g., the likelihood ratio); thus $\mathbf{C}\subsetneq\Gamma\subsetneq\mathbf{U}$. On the other hand, if we allowed $f(b)$ to be all functions of $b$, not just increasing functions, then $\Gamma=\mathbf{U}$.
@@ 2.2 Notation, assumptions, and further background
We will also assume that there exists some $b$ such that $L_{b}(0,1)>1$. Otherwise a test family may always reject with probability 1 and remain type-I risk safe, making the problem uninteresting. This assumption is akin to not studying $\alpha=1$ in the traditional Neyman-Pearson paradigm, which is uninteresting for the same reason. We also make the necessary (quite modest) measure-theoretic assumptions so that (2.3) holds. See Appendix B.1 for precise details. Finally, going forward, we assume that we are testing a point null $\mathcal{P}=\{P\}$ against a point alternative $\mathcal{Q}=\{Q\}$.
We end this section by noting how one might modify the notation we use throughout the paper.
Remark 2.10 (Koning representation of post-hoc hypothesis tests).
For the choice $L_{b}(0,1)=b$, koning2024continuous recently suggested merging the loss into the decision space of the test. Instead of a test mapping to $[0,1]$, he rescales the range to $0,L_{b}(0,1)]\supseteq[0,1]$. One could do the same thing throughout this paper. Such a transformation amounts to rescaling our usual test family $\delta$ and defining a new test family $f$ given by $f(X,b)=L_{b}(0,1)\delta(X,b)$. In this case, type-I risk safety is the condition $\mathbb{E}_{P}[\sup_{b}f(X,b)]\leq 1$. That is, the test itself becomes an e-value, as opposed to the quantity $\sup_{b}L_{b}(0,\delta(X,b))$. While Koning’s representation definitely has a strong appeal, in this paper we opt to work with the standard range of $[0,1]$ in the hope that it makes the framework more familiar to those new to post-hoc testing and e-values.
@@ 3 Warmup: Likelihood ratio tests
In the case of a single loss function, the Neyman-Pearson lemma (neyman1933ix) implies that the likelihood ratio test is most powerful. What can we say about likelihood ratio-style tests in a setting with multiple losses?
@@
$\mathbb{E}_{P}[\phi^{\textsc{np}}(X,b)]=L_{b}^{-1}(0,1).$
We use the superscript “NP” to refer to Neyman-Pearson. Note that the second term $\gamma\mathbf{1}\left\{\Lambda(X)=\kappa(b)\right\}$ is only relevant for discrete distributions. We refer to shao2008mathematical for a modern treatment of the Neyman-Pearson lemma and the optimality of the likelihood ratio test.
Let $\delta$ be an arbitrary $\Gamma$-admissible test family and suppose that it acts as $\phi^{\textsc{np}}(\cdot,b^{*})$ on some $b^{*}$ (that is, $\delta(X,b^{*})=\phi^{\textsc{np}}(X,b^{*})$). How does it behave for $b\neq b^{*}$? Naively, one would like to set $\delta(\cdot,b)=\phi^{\textsc{np}}(\cdot,b)$ on each $b$. Such a test family would not be type-I risk safe, however. This is easy to see, even in the case of two losses. Suppose that $\delta(X,b_{1})=\phi^{\textsc{np}}(X,b_{1})$ and $\delta(X,b_{2})=\phi^{\textsc{np}}(X,b_{2})$ with $b_{1}<b_{2}$. Since $L_{b_{2}}(0,1)>L_{b_{1}}(0,1)$ we have $\kappa({b_{2}})>\kappa({b_{1}})$. If an adversary $B$ plays loss $b_{2}$ for $X$ with $\Lambda(X)\geq\kappa({b_{2}})$ and $b_{1}$ for $X$ with $\kappa({b_{1}})\leq\Lambda(X)<\kappa({b_{2}})$, then $\mathbb{E}_{P}[L_{B}(0,\delta(X,B))]=L_{b_{2}}(0,1)P(\Lambda(X)\geq\kappa({b_{2}}))+L_{b_{1}}(0,1)P(\kappa(b_{1})\leq\Lambda(X)<\kappa({b_{2}}))>1$, so $\mathcal{R}_{P}(\delta)>1$.
However, for $b\leq b^{*}$, $\delta(\cdot,b)$ can act as $\phi^{\textsc{np}}(\cdot,b^{*})$ and remain type-I risk safe. (In fact, in the discrete case, for those $X$ such that $\Lambda(X)=\kappa(b)$ it can even raise the probability of rejection due to the extra freedom afforded by the fact that $L_{b}(0,1)\leq L_{b^{*}}(0,1)$.) For $b>b^{*}$ meanwhile, $\delta(\cdot,b)$ must lower the probability of rejection, but may still reject. That is, even in the continuous case, in contrast to the NP setting with a single loss function, if we allow randomized test families then admissible test families will always be randomized. Formally, Proposition 3.1 demonstrates that, for any $\Gamma\supseteq\mathbf{C}$, the only admissible test family which plays as $\phi^{\textsc{np}}(\cdot,b^{*})$ on $b^{*}$ has the following form:
@@
This representation is perhaps more intuitive to digest. If an adversary $B$ plays $B(X)=b^{*}$ for all $X$, the expected type-I risk $\mathbb{E}_{P}[L_{B}(1,\delta(X,B)]$ is precisely 1 by definition of the likelihood ratio test. For $b\leq b^{*}$ we have $L_{b}(0,1)\leq L_{b^{*}}(0,1)$ so an adversary playing such a $b$ can only lower the type-I risk and $\delta$ can play as $\phi^{\textsc{np}}(X,b^{*})$. For $b>b^{*}$, $\delta$ may still reject, but it must lower its rejection probability to make up for the increase in type-I risk. We set $\delta$ such that $L_{b}(0,1)\delta(X,b)=L_{b^{*}}(0,1)\phi^{\textsc{np}}(X,b^{*})$ so that the type-I risk stays the same. We note that koning2023markov also studied (3.3) (see his example 9). He cites it as an example of a post-hoc test, though he does not prove that it is admissible.